OBSERVATIONS OF THE C_{6}H^{−} ANION IN IRC+10216

Y. KASAI, National Institute of Information and Communications Technology (NICT), Koganei, Tokyo 184-8795, Japan; E. KAGI, Faculty of Information Science, Hiroshima City University, Asaminami-ku, Hiroshima 731-3194, Japan; \underline{K. KAWAGUCHI}, Department of Chemistry, Faculty of Science, Okayama University, Okayama 700-8530, Japan.

Existence of negative ions in space has been discussed for a long time. Although 13 positive ions have been known in interstellar space, the first identification of the negative ion C_{6}H^{−} was carried out quite recently by McCarthy et al.^{a}. In the present study, radio astronomical data of C_{6}H^{−} observed in a late tyep star IRC+10216 with Nobeyama 45-m and IRAM 30-m telescopes have been analyzed by the local thermodynamic equilibrium approximation to give the column density of 6.1-8.0 x 10^{12} cm^{−2} and the excitation temperature of 32 ± 3 K with an assumed source size of 30” ± 3”. The abundance of C_{6}H^{−} was estimated to be 3.4 of C_{6}H^{−} and C_{6}H indicate that the distribution of C_{6}H^{−} is in inner region of the circumstellar envelope than C_{6}H. The production mechanism of C_{6}H^{−} is discussed.

aM. C. McCarthy, C. A. Gottlieb, H. Gupta, P. Thaddeus, Astrophys. J. \underline{652}, L141 2006